OpenStudy (anonymous):

How would you solve this using quadratic formula: 5x^2=25x

OpenStudy (saifoo.khan):

You can use a shortcut as well. Do you want to show the working as well>

OpenStudy (anonymous):

Use the quadratic formula, while showing work, yes.

OpenStudy (saifoo.khan):

Ok, 5x^2 - 25x = 0 Where the value of: a = 5 b = -25 c = 0 Use these values in quadratic formula.

OpenStudy (anonymous):

Ok. Thanks.

OpenStudy (anonymous):

5x^2=25x 5x^2 - 25x = 0 5x ( x - 5 ) = 0 ->x = 0, x = 5